A Pedialyte solution contains Na+ 41 mEq/L, K+ 21 mEq/L, Cl− 33 mEq/L, citrate3− 29 mEq/L, and glucose (C6H12O6) 25 g/L.

1.  What is the total positive charge of the electrolytes?

2.   What is the total negative charge of the electrolytes?
Answer #1

Part 1 :

Total positive charge is mainly due to total charge of cations = Na+ + K+ = 41 + 21 = 62 mEq/L

Part 2:

Total negative charge is mainly due to total charge of anions = Cl- + Citrate -3 = 33 +29 = 62 mEq/L